OpenStudy (anonymous):

The height in feet of an object dropped from an 18- foot stepladder is given by h(t)=-16t^2+18 , where t represents the time in seconds after the object is dropped. How long does it take the object to hit the ground? (Hint: The height is 0 when the object hits the ground. Round to the nearest hundredth of a second.) Please show steps.

OpenStudy (anonymous):

0=-16t^2+18 t^2=18/16 t=sqrt(18/16)

OpenStudy (anonymous):

The answer is 1.06 seconds. I am not sure how to get to the answer.

OpenStudy (anonymous):

sqrt(18/16) = 1.06 , just set the equation to equal 0. after all you are trying to solve t when height is 0 as this is how much time has passed. just do what i did.

OpenStudy (anonymous):

Thank you for the help.

OpenStudy (anonymous):

all g

OpenStudy (anonymous):

Yep.